News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 19 June 2000 8.30am EDT

During the last week an anomaly was encountered following an HETG retraction using the B-side retract limit switch.

An observation of Serpens X-1 was completed and the HETG retracted as commanded by the daily load. The B-side retract limit switch changed state from "not-retracted" to "retracted" but finished in the "not-retracted" state (at 11.38pm EDT on 6/13) indicating a bounce from the bumper spring. The LETG failed to insert for a subsequent observation of 3C273 at 12:43am EDT on 6/14 - as would be expected given the state of the HETG retract flag ("not-retracted").

Assessment of the dump data indicates that all symptoms are the same as the A-side failure of 8 May 2000, i.e., the switch appears to function, but the trigger point has moved. The HETG was inserted manually to continue with science observations for the rest of the week and will be retracted manually at 10.30pm EDT on 6/19 after completing a scheduled HETG observation. The initial manual commanding at 9:40pm EDT on Jun 14 inserted the HETG but the insert limit switch changed to the not-inserted state at the completion of the move. This action was not planned however is consistent with the design when the lockout-override is disabled. The HETG was commanded to the fully inserted position during a real-time pass at 8.20pm EDT on Jun 15 and the insert limit switch functioned as expected.

Two observations were impacted by the anomaly, 3C273 with ACIS-S/LETG and ACIS-S/HETG. These will be rescheduled in a later load. Alternate approaches for HETG operation from the daily load are in work and an initial schedule shows resuming observations with HETG the week of July 16. Approaches for resuming LETG observations are being assessed. A meeting is planned at the OCC to discuss in detail the grating system operations.

Other real-time procedures conducted last week included a routine CPE warm start count dump on 6/12/00 (no value change), and the sixth ACIS squeegee mode test run at 2pm EDT on 6/15/00. Initial assessment of the Squeegee mode results shows good agreement with the lab results.

The planned observations for the coming week are as follows.

------------------------------------------
GX 301-2               ACIS-S/HETG     Jun 19 (manual retract)
NGC 4395               ACIS-S
M51                    ACIS-S          Jun 20
Radiation Belts
M31 (Series)           HRC-I
0519-69.0              ACIS-S          Jun 21
PSR B0540-69           ACIS-S
PSR B0540-69 (Series)  HRC-I
PSR B0540-69 (Series)  HRC-S           Jun 22
MS 1006.0+1202         ACIS-S
Ophiuchi Mosaic #1     ACIS-I
CEN X-4                ACIS-S
Radiation Belts                        Jun 23
A1413                  ACIS-I
BD Cluster 1338        ACIS-I
RBS 1223               ACIS-S          Jun 24
NGC 4736               ACIS-S
ARP 220                ACIS-S
3C 324                 ACIS-S          Jun 25
Radiation Belts

------------------------------------------

All spacecraft subsystems, with the exception of HETG as above, continued to operate nominally.

Return to Status Reports